Grade 12Algebra

the no. of ways of selecting 8 balls out of an infinite no. of white, red, yelow and brown balls is:

ans: 11C3

basit ali
14 Years agoGrade 12

4 Answers

Ashwin Muralidharan IIT Madras
14 Years ago

HI Basit,

 

Say we select x1 white balls, x2 red and x3 yello balls from the set of infinite balls.

 

Then x1+x2+x3=8 where each Xi ≥ 0.

The number of non-negative integral solutions of the eqn x1+x2+...xn = r is n+r-1Cr (This is dervided from Binomial Co-eff method).

So the solution of above eqn is 11C8 = 11C3.

 

Hope that helps.

 

Best Regards,

Ashiwn (IIT Madras).

Ashwin Muralidharan IIT Madras
14 Years ago

X1+X2+X3+X4 = 8.

So 8+4-1C8 = 11C8.

 

Best Regards,

Ashwin (IIT Madras).

basit ali
14 Years ago

Hello , Ashwin 

  How there is 11C8=11C3   ( PLZ EXPLAIN IN BRIEF.........)

      THANKS!

Ashwin Muralidharan IIT Madras
14 Years ago

Hi Basit,

 

That is because nCr = nCn-r

Which is one of the most importatn property of Combinatorics.... The Proof is quite easy.

 

Use the factorial notation, you will see nCr = n!/{r!(n-r)!}

and nCn-r = n!/{(n-r}!r!}..... Clearly both are same.

 

So 11C3 = 11C8.

 

Best Regards,

Ashwin (IIT Madtras).